# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

### v2

@@ Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. Furthermore, a major limitation of prior approaches is the lack of temporal modeling to capture time-dependent relationships among dynamically evolving entities in a scene. To address these challenges, we propose Temporal Equivariant Scene Graph Neural Network (TESGNN), consisting of two key components: (1) an Equivariant Scene Graph Neural Network (ESGNN), which extracts information from 3D point clouds to generate scene graph while preserving crucial symmetry properties, and (2) a Temporal Graph Matching Network, which fuses scene graphs generated by ESGNN across multiple time sequences into a unified global representation using an approximate graph-matching algorithm. Our combined architecture TESGNN outperforms current state-of-the-art methods in scene graph generation, achieving higher accuracy and faster training convergence. Moreover, we show that leveraging the symmetry-preserving property produces a more stable and accurate global scene representation compared to existing approaches. Last but not least, it is computationally efficient and easily implementable using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges. Our source code is publicly available at: https://github.com/HySonLab/TESGraph.
I Introduction
@@
Figure 1: A visualization of multi-view scene graph from multiple 3D point cloud sequences. Our proposed TESGNN first generates local scene graphs for each sequence using Equivariant GNN. Then, the local scene graphs are merged by passing through a temporal layer to form a global scene graph representing the entire scene.
@@
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems [1, 2, 3]. The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach [4], it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques [5] provide richer spatial information but are resource intensive and particularly susceptible to noise interference [6]. Recent innovations, such as LMNav [7] and VLMap [8], leverage Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques [9]. Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work [10, 11], researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. The latest advancements [12] propose hybrid approaches that integrate both 2D and 3D data to further refine scene graph representations. However, based on our analysis, existing approaches reveal two major limitations.
@@
Current methods struggle with noisy, multi-view 3D data, leading to inconsistent scene graphs due to sensitivity to camera angles and poor symmetry preservation. Enhancing Graph Neural Networks (GNNs) with invariant node and edge embeddings and using Equivariant GNNs [13, 14] can address this, though prior works [11, 12] haven’t explored these approaches. Additionally, existing methods lack robust strategies for building coherent global scene graphs from multiple point cloud sequences. For example, [11] uses heuristic averaging to merge local graphs but skips optimization, resulting in global graphs that lack semantic coherence and structural consistency.
@@
  • Equivariant Scene Graph Neural Network (ESGNN) for Scene Graph Extraction: extracts scene graphs from 3D point clouds using an Equivariant GNN with Equivariant Graph Convolution and Feature-wise Attention layers, ensuring rotational and translational invariance. This design delivers high-quality scene representations with superior accuracy, fewer training iterations, and reduced computational demands compared to existing methods—marking the first application of symmetry-preserving Equivariant GNNs for this task.
  • Temporal Scene Graph Matching for Global Scene Representation: fuses the local scene graphs into a unified global representation by solving a graph matching problem based on node embedding similarity, with ESGNN as the backbone scene graph extractor. This embedding-based temporal approach achieves robust multi-view scene understanding, outperforming other extractors and enabling applications like multi-agent coordination and navigation in dynamic environments.
II Related Work
@@
Graph representation learning encodes structural and relational data into low-dimensional embeddings using methods like GNNs, Graph Transformers [15, 16, 17], and Graph Attention Networks [18], enabling tasks such as scene graph understanding without dense 3D data. For 3D applications, standard GNNs lack rotational symmetry; recent SE(3)/E(3)-equivariant architectures like EGNN [19] and Equiformer [14] address this by preserving geometric symmetries during transformations, critical for point clouds [20], molecular structures, and spatial reasoning.
Scence Graph Understanding
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as mobile robot navigation. Originally introduced as a structure to represent object instances and their relationships within a scene [21], scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images [21, 22], videos [23], and point clouds [10]. [10] pioneered 3DSSG, the first method for generating 3D scene graphs from point clouds, accompanied by the annotated 3RScan dataset. Building on this, [11] introduced Scene Graph Fusion, which employs feature-wise attention to improve scene graph representations, followed by an incremental approach [12] that integrates both RGB image sequences and sparse 3D point clouds. Scene Graph Fusion remains the state-of-the-art method for scene graph generation from 3D point clouds. Our approach builds upon this by integrating Feature-wise Attention to optimize the Message-Passing process, thereby enhancing the quality of the generated scene graphs.
@@
Temporal graph learning has gained significant attention for modeling dynamic relationships over time, with applications in traffic prediction [24, 25, 26]. Traditional GNNs assume static graph structures, limiting their applicability in scenarios like video analysis and multi-view scene understanding. To address this, Temporal GNNs [27, 28] incorporate time-series learning to capture evolving patterns. In scene graph matching, methods such as SGAligner [29] and SG-PGM [30] align nodes across graphs but rely on ground truth graphs, making them impractical for real-world tasks where predicted graphs contain noise, ambiguous edges, and node permutations. To overcome these challenges, we introduce a symmetry-preserving temporal layer in ESGNN, leveraging equivariant properties to merge scene graphs across time steps. Unlike prior approaches that treat temporal information as sequential snapshots, our method integrates graph matching to construct a unified representation that maintains temporal consistency while reducing computational overhead.
Figure 2: Overview of our proposed TESGNN. Our approach takes sequences of point clouds a) as input to generate a geometric segmentations b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
@@
Reconstructing large-scale point clouds is inherently challenging due to their irregular structure, often comprising multiple sub-point cloud sequences [31]. To tackle this issue, we introduce a temporal model that merges graphs generated from these sequences by utilizing embeddings from our proposed models. This approach not only improves integration across sequences but also extends to applications such as multi-agent SLAM and dynamic environment adaptation [32, 33], offering an efficient, lightweight solution based on multi-view scene graph representations.
III Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph from sequence of point clouds. The framework consists of three key phases: Feature Extraction (a-c), discussed in Section III-B; Scene Graph Extraction (d-f); and Temporal Graph Matching (g-h), detailed in Section III-C. Our main works, scene graph extraction and temporal graph matching, are further elaborated in Sections IV and V.
III-A Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section III-B.
@@ III-A Problem Formulation
[…]
@@
Let us say we have a global point cloud, that is,
@@
where $\Phi_{graph}$ is the model matching the embedding vectors together, which will be described in the later section.
III-B Feature Extraction
In this phase, the framework extracts the feature for scene graph generation, following the two main steps:
@@ III-B Feature Extraction
    Point Cloud Reconstruction: The proposed framework takes the point cloud data, which can be reconstructed from various techniques such as ORB-SLAM3 or HybVIO [34] as the input. However, for the objective validation purpose of scene graph generation, we use the indoor point cloud dataset 3RScan [35] for ground truth data $D_{i}$.
@@
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet [36] $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section IV.
III-C Scene Graph Extraction and Temporal Graph Matching
In this phase, the framework processes the input from feature extraction (Section III-B) to generate the scene graph:
  • Properties and Neighbor Graph Extraction: From the point cloud, we extract features including the centroid $\overline{\mathbf{p}}_{i}\in\mathbb{R}^{3}$, standard deviation $\boldsymbol{\sigma}_{i}\in\mathbb{R}^{3}$, bounding box size $\mathbf{b}_{i}\in\mathbb{R}^{3}$, maximum length $l_{i}\in\mathbb{R}$, and volume $\nu_{i}\in\mathbb{R}$. We create edges between nodes only if their bounding boxes are within 0.5 meters of each other, following [11].
  • Scene Graph Extraction and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section IV) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section V). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. ESGNN is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in [10]. Meanwhile, our Temporal Graph Matching is trained to minimize the representation distance between identical nodes with contrastive loss [37].
@@ III-C Scene Graph Extraction and Temporal Graph Matching
IV Equivariant Scene Graph Neural Network (ESGNN)
To build a network architecture that effectively generate scene graph from point cloud, we propose a combination of Equivariant Graph Convolution Layers [19] and the Graph Convolution Layers with Feature-wise Attention [11].
IV-A Graph Initialization

### v3

@@ Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. Furthermore, a major limitation of prior approaches is the lack of temporal modeling to capture time-dependent relationships among dynamically evolving entities in a scene. To address these challenges, we propose Temporal Equivariant Scene Graph Neural Network (TESGNN), consisting of two key components: (1) an Equivariant Scene Graph Neural Network (ESGNN), which extracts information from 3D point clouds to generate scene graph while preserving crucial symmetry properties, and (2) a Temporal Graph Matching Network, which fuses scene graphs generated by ESGNN across multiple time sequences into a unified global representation using an approximate graph-matching algorithm. Our combined architecture TESGNN shown to be effective compared to existing methods in scene graph generation, achieving higher accuracy and faster training convergence. Moreover, we show that leveraging the symmetry-preserving property produces a more stable and accurate global scene representation compared to existing approaches. Finally, it is computationally efficient and easily implementable using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges.
1 Introduction
@@
Figure 1: A visualization of a multi-view scene graph from multiple 3D point cloud sequences. Our proposed TESGNN first generates local scene graphs for each sequence using Equivariant GNN. Then, the local scene graphs are merged by passing through a temporal layer to form a global scene graph representing the entire scene.
@@
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems Kim et al. (2024); Li et al. (2022b); Koch et al. (2024a). The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach Pfaff et al. (2006), it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques Jin et al. (2024) provide richer spatial information but are resource intensive and particularly susceptible to noise interference Li et al. (2022a). Recent innovations Shah et al. (2022); Huang et al. (2023); Yamazaki et al. (2024) leverage Vision Language Models (VLMs) and Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques Chang et al. (2023). Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work Wald et al. (2020); Wu et al. (2021b), researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. Latest advancements Wu et al. (2023); Koch et al. (2024b) propose hybrid, multi-modal approaches that integrate both 2D and 3D data to further refine scene graph representations. These hybrid approaches often yield better prediction results by combining image encoders or VLMs with GNNs for scene graph generation.
@@
However, based on our analysis, existing GNN architectures for scene graph generation have several limitations. Current methods struggle with noisy, multi-view 3D data, leading to inconsistent scene graphs due to sensitivity to camera angles and poor symmetry preservation. Enhancing GNNs with invariant node and edge embeddings and using Equivariant GNNs Köhler et al. (2020); Liao & Smidt (2023) can address this, though prior works Wu et al. (2021b; 2023) haven’t explored these approaches. Additionally, existing methods lack robust strategies for building coherent global scene graphs from multiple point cloud sequences. For example, Wu et al. (2021b) uses heuristic averaging to merge local graphs but skips optimization, which could potentially result in global graphs that lack semantic coherence and structural consistency.
@@
  • Equivariant Scene Graph Neural Network (ESGNN) for Scene Graph Extraction: extracts scene graphs from 3D point clouds using an Equivariant GNN with Equivariant Graph Convolution and Feature-wise Attention layers, ensuring rotational and translational invariance. This design delivers high-quality scene representations with superior accuracy, fewer training iterations, and reduced computational demands compared to existing methods - marking the first application of symmetry-preserving Equivariant GNNs for this task.
  • Temporal Scene Graph Matching for Global Scene Representation: fuses the local scene graphs into a unified global representation by solving a graph matching problem based on node embedding similarity, with ESGNN as the backbone scene graph extractor. This embedding-based approach is robust for multi-view scene understanding, being more effective compared to other methods and enabling applications in multi-agent coordination and navigation in dynamic environments.
2 Related Work
@@
Graph representation learning encodes structural and relational data into low-dimensional embeddings using methods like GNNs, Graph Transformers Yun et al. (2019); Kim et al. (2022); Cai et al. (2023), and Graph Attention Networks Veličković et al. (2018), enabling tasks such as scene graph understanding without dense 3D data. For 3D applications, standard GNNs lack rotational symmetry; recent SE(3)/E(3)-equivariant architectures like EGNN Satorras et al. (2021) and Equiformer Liao & Smidt (2023) address this by preserving geometric symmetries during transformations, critical for point clouds Uy et al. (2019), molecular structures, and spatial reasoning.
Scene Graph Understanding
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as robot navigation Honerkamp et al. (2024); Pham et al. (2025). Originally introduced as a structure to represent object instances and their relationships within a scene Johnson et al. (2015), scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images Johnson et al. (2015); Armeni et al. (2019) and videos Qi et al. (2018). 3DSSG Wald et al. (2020) pioneered scene graph generation from 3D point clouds. Building on this, SGFN Wu et al. (2021b) employs Feature-wise Attention to improve scene graph representations, followed by an incremental approach Wu et al. (2023) that integrates both RGB image sequences and sparse 3D point clouds. To the best of our knowledge, SGFN remains state-of-the-art for scene graph generation from 3D point clouds, without incorporating image encoder or VLMs. Our approach adopts this Feature-wise Attention to optimize the Message-Passing process and enhance the generated scene graphs. Additionally, it is important to note that while recent works Wu et al. (2023); Koch et al. (2024b); Saxena et al. (2025) focus on utilizing vision–language encoders, our work investigates a different problem: How to improve the GNN backbone design for 3D understanding, leading to better generalization and faster convergence.
@@
Temporal graph learning has gained significant attention for modeling dynamic relationships over time, with applications in traffic prediction Yu et al. (2018); Li et al. (2018); Nguyen et al. (2024a). Traditional GNNs assume static graph structures, limiting their applicability in scenarios like video analysis and multi-view scene understanding. To address this, Temporal GNNs Rossi et al. (2020; 2023) incorporate time-series learning to capture evolving patterns. In scene graph matching, methods such as SGAligner Sarkar et al. (2023) and SG-PGM Xie et al. (2024) align nodes across graphs but rely on ground truth graphs, making them impractical for real-world tasks where predicted graphs contain noise, ambiguous edges, and node permutations. To overcome these challenges, we introduce a symmetry-preserving temporal layer in ESGNN, leveraging equivariant properties to merge scene graphs across time steps. Unlike prior approaches that treat temporal information as sequential snapshots, our method integrates graph matching to construct a unified representation.
Figure 2: Overview of our proposed TESGNN. Our approach takes sequences of point clouds a) as input to generate a geometric segmentation b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
@@
Reconstructing large-scale point clouds is inherently challenging due to their irregular structure, often comprising multiple sub-point cloud sequences Huang et al. (2022). To tackle this issue, we introduce a temporal model that merges graphs generated from these sequences by utilizing embeddings from our proposed models. This approach not only improves integration across sequences but also extends to applications such as multi-agent SLAM and dynamic environment adaptation Jiang et al. (2019); Zou et al. (2019), offering an efficient, lightweight solution based on multi-view scene graph representations.
3 Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph from a sequence of point clouds. The framework consists of three key phases: Feature Extraction (a-c), discussed in Section 3.2; Scene Graph Extraction (d-f); and Temporal Graph Matching (g-h), detailed in Section 3.3. Our main works, scene graph extraction and temporal graph matching, are further elaborated in Sections 4 and 5.
3.1 Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section 3.2.
@@ 3.1 Problem Formulation
The Semantic Scene Graph is denoted as $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ represent sets of entity nodes and directed edges, respectively. In this case, each node $v_{i}\in\mathcal{V}$ contains an entity label $l_{i}\in L$, point clouds $\mathcal{P}_{i}$, an 3D Oriented Bounding Box (OBB) $b_{i}$, and a node category $c_{i}^{\text{node}}\in\mathcal{C}^{\text{node}}$. Conversely, each edge $e_{i\rightarrow j}\in\mathcal{E}$, connecting node $v_{i}$ to $v_{j}$ where $i\neq j$, is characterized by an edge category or semantic relationship denoted by $c_{i\rightarrow j}^{\text{edge}}\in\mathcal{C}^{\text{edge}}$, or can be written in a relation triplet $\left<\textit{subject, predicate, object}\right$>. Here, $L$, $\mathcal{C}^{\text{node}}$, and $\mathcal{C}^{\text{edge}}$ represent the sets of all entity labels, node categories, and edge categories, respectively. Given the 3D scene data $D_{i}$ and $D_{j}$ that represent the same point cloud of a scene but from different views (rotation and transition), we try to predict the probability distribution of the equivariant scene graph prediction in which the equivariance is preserved:
@@
We denote a global point cloud, that is,
@@
where $\Phi_{graph}$ is the model matching the embedding vectors together, which will be described in a later section.
3.2 Feature Extraction
In this phase, the framework extracts the features for scene graph generation, following the two main steps:
@@ 3.2 Feature Extraction
    Point Cloud Reconstruction: The proposed framework takes the point cloud data, which can be reconstructed from various techniques such as ORB-SLAM3 or HybVIO Campos et al. (2021) as the input. However, for the objective validation purpose of scene graph generation, we use the indoor point cloud dataset 3RScan Wald et al. (2019) for ground truth data $D_{i}$.
@@
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet Charles et al. (2017) $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section 4.
3.3 Scene Graph Extraction and Temporal Graph Matching
In this phase, the framework processes the input from feature extraction (Section 3.2) to generate the scene graph:
  • Properties and Neighbor Graph Extraction: From the point cloud, we extract features including the centroid $\overline{\mathbf{p}}_{i}\in\mathbb{R}^{3}$, standard deviation $\bm{\sigma}_{i}\in\mathbb{R}^{3}$, bounding box size $\mathbf{b}_{i}\in\mathbb{R}^{3}$, maximum length $l_{i}\in\mathbb{R}$, and volume $\nu_{i}\in\mathbb{R}$. We create edges between nodes only if their bounding boxes are within 0.5 meters of each other, following Wu et al. (2021b).
  • Scene Graph Extraction and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section 4) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section 5). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. ESGNN is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in Wald et al. (2020). Meanwhile, our Temporal Graph Matching is trained to minimize the representation distance between identical nodes with contrastive loss Hadsell et al. (2006).
4 Equivariant Scene Graph Neural Network (ESGNN)
@@
To build a network architecture that effectively generate scene graph from point cloud, we propose a combination of Equivariant Graph Convolution Layers Satorras et al. (2021) and the Graph Convolution Layers with Feature-wise Attention Wu et al. (2021b).
4.1 Graph Initialization